### v3

@@ 3.7 Modality Ablation Study
Integrating both modalities through CMAFNet yields consistent and substantial improvements over either single-modality baseline. The fused configuration achieves mAP50 of 0.244, representing gains of 26.4% over RGB-only and 61.6% over depth-only settings. More notably, APs increases from 0.073 to 0.089, demonstrating a 21.9% relative improvement in small-object detection—the predominant challenge in transmission-line inspection where 94.5% of defect instances qualify as small objects. The recall metrics exhibit similar trends: insulator contamination recall rises from 0.367 (RGB) to 0.477 (RGB+Depth), and conductor strand breakage recall improves from 0.310 to 0.365. These category-specific gains indicate that depth information provides complementary boundary and surface discontinuity cues that enhance detection of geometrically subtle defects exhibiting low chromatic contrast against cluttered backgrounds.
The precision-recall trade-off merits attention. While RGB-only achieves the highest precision (0.677), it exhibits low recall (0.187), indicating a conservative detection strategy that misses numerous true positives. The fused model accepts a moderate precision reduction (0.452) in exchange for substantially higher recall (0.232), yielding a net improvement in overall detection performance as reflected in mAP metrics. This trade-off aligns with practical inspection requirements where false negatives carry greater risk than false positives.
From a computational perspective, the multimodal configuration introduces marginal overhead—5.5M parameters and 12.9 GFLOPs compared to 4.9M and 12.4 GFLOPs for single-modality baselines—while maintaining real-time inference at 222.3 FPS. These results collectively validate both the effectiveness of depth as a complementary modality and the efficiency of the proposed cross-modal alignment and fusion strategy.

### v4

@@ 3.7 Modality Ablation Study
Integrating both modalities through CMAFNet yields consistent and substantial improvements over either single-modality baseline. The fused configuration achieves mAP50 of 0.244, representing gains of 26.4% over the RGB-only setting and 61.6% over the depth-only setting. More notably, APs increases from 0.073 to 0.089, demonstrating a 21.9% relative improvement in small-object detection—the predominant challenge in transmission-line inspection where 94.5% of defect instances qualify as small objects. The recall metrics exhibit similar trends: insulator contamination recall rises from 0.367 (RGB) to 0.477 (RGB+Depth), and damaged pole recall improves from 0.310 to 0.365. These category-specific gains indicate that depth information provides complementary boundary and surface discontinuity cues that enhance the detection of geometrically subtle defects exhibiting low chromatic contrast against cluttered backgrounds.
The precision-recall trade-off merits attention. While the RGB-only configuration achieves the highest precision (0.677), it exhibits low recall (0.187), indicating a conservative detection strategy that misses numerous true positives. The fused model accepts a moderate precision reduction (0.452) in exchange for substantially higher recall (0.232), yielding a net improvement in overall detection performance as reflected in mAP metrics. This trade-off aligns with practical inspection requirements where false negatives carry greater risk than false positives.
From a computational perspective, the multimodal configuration introduces marginal overhead—5.5M parameters and 12.9 GFLOPs compared to 4.9M and 12.4 GFLOPs for the single-modality baselines—while maintaining real-time inference at 222.3 FPS. These results collectively validate both the effectiveness of depth as a complementary modality and the efficiency of the proposed cross-modal alignment and fusion strategy.